In a chemical manufacturing plant, a certain type of chemical B is produced

through reactions in chemical A. Through observation the officers on duty have

noted that the rate of conversion from chemical A to B is proportion to the

amount of chemical A present at any time, further investigation has revealed

that 10% amount has been converted in first five minutes.

a) The plant manager wishes to know the percentage of chemical A that will

be converted in 20 minutes.

b) The plant manager wishes to know the time it will take to convert 60

minutes of chemical A.


1
Expert's answer
2021-05-10T13:07:01-0400

For this

We first calculate t1/2 that is half conversion time


"{N(t)\over N_o} = ({1\over 2})^{t\over t_{1\over 2}}" .......(1)

"N(t)=" amount at time t (let it be 100)

"N_o=" initial amount

By putting values

Given t = 5 min

Amount converted = 10%

So

"N_o=" 90

(Note we use ln to solve)

We got

t1/2=34.5


a) at t=20 and t1/2=34.5


Putting values in 1 we got

"{N(20)\over N_o}=({1\over 2})^{20\over 34.5}"

And solving

We got

N(20) = 67.3 %

Converted = 100-67.3

=32.7%


b)

So N(t)=100-60

=40%

No= 100

t1/2=34.5

Putting values in (1)


And we got

"{40\over100 }=({1\over 2})^{t\over {34.5}}"

taking ln both side , we got

t = 45.5 minutes